OpenStudy (yadanawala):

The number of bacteria in a certain population increases according to a continuous exponential growth model, with a growth rate parameter of 6.1% per hour. How many hours does it take for the size of the sample to double?.............. Note: This is a continuous exponential growth model. Do not round any intermediate computations, and round your answer to the nearest hundredth.

OpenStudy (kropot72):

Let N be the number of bacteria as function of the time in hours, and let No be the initial number of bacteria. \[\large N=N _{0}e ^{0.061t}\] \[\large \frac{N}{N _{0}}=2=e ^{0.061t}\] \[\large 0.061t=\ln 2\] \[\large t=\frac{\ln 2}{0.061}=you\ can\ calculate\]